OpenStudy (anonymous):

Divide (1/3)*[(n+1)^3/n^3] I'm supposed to find (1/3)(1+1/n)^3 but I'm finding (1/3)*(1+(1/n))^3/n^2 Thanks for your futur help :-)

OpenStudy (mathmale):

For clarity, would you please type out your expression in Equation Editor or by hand in the Draw utility, below. \[\frac{ 1 }{ 3 }[(n+1)^3/n^3] ~\rightarrow~\frac{ 1 }{ 3 }\frac{ (n+1)^3 }{ n^3 }\]

OpenStudy (mathmale):

Is division all that you have to do? If so, please expand the binomial (n+1)^3.

OpenStudy (anonymous):

 I got this. I don't get it when you tell me to expand (n+1)^3